# Closed-loop Long-horizon Robotic Planning via Equilibrium Sequence Modeling

### v4

@@ 3.4 Practical Implementation
This section discusses the implementation of the proposed equilibrium planner. To enable effective training while interacting with the environment, the following two modules are carefully devised to complement the planner: an experience memory that caches all equilibrium plans and their feedback during equilibrium solving, and a world model to estimate the closed-loop feedback in the absence of environmental interactions. The planning framework is illustrated in Fig. 3.
@@
Equilibrium Experience Memory. During the training process, our equilibrium model interacts with the environment only when the inner loop reaches the equilibrium point. This results in a small number of equilibrium points, which may not be sufficient for supervised training. To improve our training efficiency and stability, we opt to cache all previously obtained equilibrium points, along with their environmental feedback, in an experience memory. Thereafter, these equilibrium points can be sampled repeatedly for versatile training purposes. For example, for the planner, we randomly sample a batch of equilibrium points at each training epoch, which are paired with the ground truth for supervised training. In particular, the most recent equilibrium points are sampled more frequently to reduce distribution shift. Next, we describe another crucial component of our framework.
@@
Internal Feedback from World Model. Due to inefficiency of interacting with the environment, we construct a world model (Ha & Schmidhuber, 2018) to provide the necessary feedback in closed-loop planning. Our world model takes the environmental context, task instruction and current plan as inputs and predicts some basic types of feedback. This definition is slightly simpler than the commonly used world model, which requires simulation of the environmental states, and therefore may be easier to train. Concretely, we implement the world model with an LLM and finetune it on the planner’s equilibrium feedback over all iterations for better generalizability. And during inference, we alternate between using external and internal feedback in closed-loop robotic planning.
Figure 3: Illustration of our proposed framework. It incorporates (1) a memory containing all the equilibrium experience during inference, (2) a self-refining planner trained on equilibrium plans with the ground truth, and (3) a world model trained on the memory to simulate environmental feedback.
4 Experiments
4.1 Experimental Settings
Metrics. We use executability (Exec.), success rate (SR), goal conditions recall (GCR) following Hu et al. (2024). Exec. evaluates whether the plan can be executed in given the environment, SR refers to whether the goal is accomplished, and GCR measures the proportion of goal conditions achieved. To examine the closed-loop planning capabilities, we study two test setups, without error correction or with up to 10 corrections, the latter allowing for self-correction based on environmental feedback. We also evaluate the computational efficiency by measuring TFLOPS at inference time.
Baselines. Our method is compared with Tree-Planner (Hu et al., 2024), SELF-REFINE (Madaan et al., 2023), and a supervised finetuned planner. They are all reproduced using Llama 3 8B (Dubey et al., 2024). In addition, we consider several baseline methods that call the GPT-3.5 API, including ProgPrompt (Singh et al., 2023), Zero-shot Planner (Huang et al., 2022) and two self-refining planners, Local Replan (Raman et al., 2022; Guo et al., 2023) and Global Replan (Shinn et al., 2023). Their results are presented for reference only. See Section B.2 for more details.
Implementation Details. Our implementation is consistent with the baseline methods by finetuning from Llama 3 8B (Dubey et al., 2024) on the VirtualHome-Env training set (paired with the equilibrium points). The number of finetuning epochs is set to 6, and the learning rate is set to 0.0002. The world model is finetuned on all planner interactions for 5 epochs using the same learning rate. A greedy LLM sampling strategy is used in later refinement steps until convergence. Moreover, we implement the KV cache to speed up inference. Further details are provided in Section B.3.
@@ 4.1 Experimental Settings
Table 1: Performance on VirtualHome-Env without correction. Our planner achieves state-of-the-art performance in most evaluations. Note that the Exec. metrics are marked in gray because they are already high and can easily exceed 99% with simple automated rules (by truncating illegal output).
@@
Table 2: Performance on VirtualHome-Env with up to 10 corrections. Our planner consistently leads in SR and GCR performance. The Exec. metrics are shown in gray for the same reasons in Table 1.
@@
4.2 Main Results
@@
Task fail, trackback
@@
Figure 4 further compares our self-correction process with the baseline methods. As can be seen, our method is better at incorporating environmental feedback to improve the plan, while the baselines fail by simply repeating the previous plan, or by making only local adjustments that are insufficient. This exemplifies the effectiveness of equilibrium sequence modeling for self-correction.
Table 3: Effectiveness of different types of feedback. They are measured under the constraint of up to 10 rounds of internal or external feedback. The last two rows are reported in the main tables.
@@
(a) Scaling w.r.t. inference computation
(b) Performance for long-horizon tasks
@@
(a) Effectiveness of reusing equilibrium
(b) Dynamic computation allocation
Figure 6: Ablation study on computational efficiency. Our nested equilibrium solving process reuses previous equilibrium solutions and dynamically allocates the inference budget for better efficiency. The latter is displayed as the number of fixed-point iterations with mean and standard deviation.
4.3 Ablation Study
Effectiveness of equilibrium sequence modeling is quantitatively verified in Table 2. Compared to alternative methods, our approach shows more than 11% improvement over the prompting-based self-refinement method SELF-REFINE (Madaan et al., 2023) and more than 12% improvement over the system 2-based Tree-Planner (Hu et al., 2024). Meanwhile, we maintain a simplistic supervised finetuning fashion similar to the compared methods, without intricate reinforcement learning. This validates the effectiveness of equilibrium sequence modeling in our robot task planning problem.
Scaling of performance. Here, we follow Brown et al. (2024); Snell et al. (2024); Wu et al. (2024b); OpenAI (2024) in considering the scaling w.r.t. inference computation. The results in Fig. 5(a) show that our method achieves better performance-computation tradeoff along with leading scaling w.r.t. inference computation. Thus, more inference budget can be allocated to improve its performance. Furthermore, in Fig. 5(b), we show that its performance advantage is largely due to better long-horizon planning capabilities, achieving more than twice the success rate for extremely long plans.
Computational efficiency. Although our training is slower than the baselines (36h vs. 12h) due to the equilibrium solving process for synthesizing training pairs ($\approx$24h), it exhibits a competitive inference efficiency. For example, our method takes 16h to evaluate, while Tree-Planner takes 24h. This can be attributed to our design of reusing equilibrium in nested equilibrium solving, As illustrates in Fig. 6(a), it accelerates the convergence of feedback iterations, achieving better performance ($>$55%) with significantly fewer interactions. Furthermore, we show in Fig. 6(b) that our method is able to dynamically allocate inference computation for planning tasks of different complexity.
@@ 5 Conclusion
This work proposes an equilibrium model-based LLM planner that is capable of self-refining plans from external and internal feedback. Unlike existing self-refinement methods based on prompting or sophisticated reinforcement learning, our proposed equilibrium sequence modeling allows simple supervised training of the self-refining planners. Moreover, it also enables the planner to efficiently incorporate environmental feedback or a world model for closed-loop planning. We implement the proposed approach on the VirtualHome-Env benchmark, and the experimental results suggest that it can dynamically allocate inference computation to achieve state-of-the-art planning performance. Code: https://github.com/Singularity0104/equilibrium-planner.
@@ A.2 Training Deep Equilibrium Models
Its proof is given in Section A.4. Due to the challenge of exactly computing the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ in the above gradient, existing work often approximate it via the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). Recently, Fung et al. (2022); Geng et al. (2021a) propose to drop the inverse Jacobian term using $A\approx I$, with the former justifying it in Theorem 3.1 with strong assumptions. In practice, this approximation has been used extensively in one-step gradient (Bolte et al., 2023; Luketina et al., 2016; Finn et al., 2017; Liu et al., 2019; Garima et al., 2020) and shown to be effective on Transformer-based LLMs (Choe et al., 2023).
@@ A.3 Transformer-based Deep Equilibrium Models
Recently, following the prevalence of Diffusion Transformers (Peebles & Xie, 2023), deep equilibrium models are extended to image generation tasks. Geng et al. (2023) propose generative equilibrium Transformers consisting of two modules, one using Transformer as the basic unit $f_{\theta}$ in the equilibrium model. Their method yields advanced one-step image generation results. (Bai & Melas-Kyriazi, 2024) replace most of the intermediate Transformer blocks with an equilibrium model, thus significantly reducing the number of parameters and memory usage for training and inference.
@@ A.4 Proof of Implicit Function Theorem
Proof of Theorem 1. To derive the loss gradient w.r.t. $\theta$, we begin by differentiating the equilibrium condition $x^{*}=f_{\theta}(x^{*},c)$ with respect to $\theta$. Applying the chain rule, we have:
@@ Appendix B Experimental Settings
Figure 8: Detailed statistics of VirtualHome-Env (Puig et al., 2018; Liao et al., 2019). The benchmark features (a) a large set of long-horizon plans with an average length of 10.8, and (b) 7 complex scenes containing more than 280 objects and more than 400 valid relations. For the sake of clarity, we have excluded the CLOSE and FACING relations, which are redundant for most planning tasks.
@@ B.1 Benchmark
Figure 7 visualizes a few examples sampled from the VirtualHome-Env benchmark. In each example, the planner is placed in an environment that spans a few indoor rooms and is given a detailed description of the environment. The description is originally in the form of a scene graph with objects as nodes and spatial relationships as edges, but for simplicity we present the planner with only the object nodes. The planner is then asked to generate a semantic action sequence based on a short task description. For instance, after receiving an instruction “turn on TV …”, the robot agent must first walk to the table and grab the remote control, and then point at the TV to turn it on. As can be seen, these planning tasks usually involve a rather complex scene setup, and the ground truth action sequences are quite long. We provide more detailed statistics for this benchmark in Fig. 8.
Compared to alternative embodied planning benchmarks, VirtualHome-Env features long time horizons as well as more diverse closed-loop feedback. For example, ALFRED (Shridhar et al., 2020) and ReALFRED (Kim et al., 2024) are two common embodied instruction following benchmarks, but their plan lengths are relatively short and can be determined by a few templates, making them unsuitable for long-horizon planning. PlanBench (Valmeekam et al., 2023) and TravelPlanner (Xie et al., 2024) are recent benchmarks designed specifically for LLM planning, but they do not provide closed-loop feedback during execution, which is an essential element of robotic planning. Therefore, we adopt VirtualHome-Env during the experiments, with further details presented below.
@@
    Program format feedback:“Your output does not conform to the required format”, meaning that the generated action sequence does not conform to the required format.
@@
    Invalid command feedback: “Your output has an invalid command: …”, meaning that the generated action sequence has an illegal command line.
@@
    Execution feedback: “Your output is executed incorrectly in the environment.”, meaning that the generated action sequence cannot be executed in the environment.
@@
    Task completion feedback: “You have not completed this task. The following objects and corresponding states do not meet the goals: … The following objects have wrong relative position: …”, meaning that the generated action sequence cannot complete the task, with more details about the task failure.
@@ B.2 Baselines
Our method is mainly compared with Tree-Planner (Hu et al., 2024) and SELF-REFINE (Madaan et al., 2023), both reproduced using Llama 3 8B Instruct (Dubey et al., 2024) in line with ours. The former traverses an action tree that is built by repeated plan sampling, while the latter relies on self-refinement. To reproduce them, we perform supervised finetuning of Llama 3 on the training split of VirtualHome-Env for the same number of epochs as our method, and then follow their original procedures for inference. For instance, Tree-Planner is reproduced with both settings $N\in\{25,50\}$ in action tree construction. The system prompts they use are similar to ours in Fig. 9.
We also report the results summarized by Hu et al. (2024) for reference. They additionally considered ProgPrompt (Singh et al., 2023), Zero-shot Planner (Huang et al., 2022) and two self-refinement planners, Local Replan (Raman et al., 2022; Guo et al., 2023) and Global Replan (Shinn et al., 2023). Since these baselines were implemented by calling the GPT-3.5 API instead of finetuning Llama 3, we report them in the novel scene and task track for a relatively fair comparison. It is worth noting that they adopted a smaller subset of actions and feedback, and differed in the curation of partial observations of the environment. Therefore, their results are for reference only.
There are several robotic planning baselines that we have not compared due to large environmental differences. For example, LLM-MCTS (Zhao et al., 2023) is a representative tree-search (Yao et al., 2023a) based planner. It followed Watch-and-help (Puig et al., 2021) to generate a dataset of simple embodied tasks (mostly object rearrangement tasks), while our work considers a more complex set of planning tasks. Alternative planners based on symbolic scene graph (Zhu et al., 2021; Rana et al., 2023), code (Liang et al., 2023; Sun et al., 2023), or PDDL (McDermott, 2000; Liu et al., 2023; Guan et al., 2023) are less flexible and difficult to implement in our environment.

### v5

@@ 3.4 Practical Implementation
This section discusses the implementation of the proposed equilibrium planner. To enable effective training while interacting with the environment, the following two modules are carefully devised to complement the planner: an experience memory that caches all equilibrium plans and their feedback during equilibrium solving, and a world model to estimate the feedback in the absence of environmental interactions. The complete planning framework is illustrated in Figure 3.
@@
Equilibrium Experience Memory. During the training process, our equilibrium model interacts with the environment only when the inner loop reaches the equilibrium point. This results in a small number of equilibrium points, which may not be sufficient for supervised training. To improve our training efficiency and stability, we opt to cache all previously obtained equilibrium points, along with their environmental feedback, in an experience memory. Thereafter, these equilibrium points can be sampled repeatedly for versatile training purposes. For example, for the planner, we randomly sample a batch of equilibrium points at each training epoch, which are paired with the ground truth for supervised training. In particular, the most recent equilibrium points are sampled more frequently to reduce distribution shift. Next, we describe another crucial component.
@@
Internal Feedback from World Model. Due to inefficiency of interacting with the environment, we construct a world model (Ha & Schmidhuber, 2018) to provide the necessary feedback in closed-loop planning. Our world model takes the environmental context, task instruction and current plan as inputs and predicts some basic types of feedback. This definition is slightly simpler than the commonly used world model, which requires simulation of the environmental states, and therefore may be easier to train. Concretely, we implement the world model with an LLM and finetune it on the planner’s equilibrium feedback over all iterations for better generalizability. And during inference, we alternate between using real and generated feedback at each iteration for a good balance between performance and efficiency.
@@
Table 1: Performance on VirtualHome-Env without correction. Our planner achieves state-of-the-art performance in most evaluations. Note that the Exec. metrics are marked in gray because they are already high and can easily exceed 99% with simple automated rules (by truncating illegal output). See Table 6 in the appendix for full comparisons with Tree-Planner.
@@
Table 2: Performance on VirtualHome-Env with up to 10 corrections. Our planner consistently leads in SR and GCR performance. In particular, the comparison with SELF-REFINE confirms the effectiveness of our new training method.
@@
4 Experiments
4.1 Experimental Settings
Metrics. We use executability (Exec.), success rate (SR), goal conditions recall (GCR) following Hu et al. (2024). Exec. evaluates whether the plan can be executed in given the environment, SR refers to whether the goal is accomplished, and GCR measures the proportion of goal conditions achieved. To examine closed-loop planning capabilities, we study two test settings, without error correction or with up to 10 corrections, the latter allowing self-correction based on environmental feedback. We also evaluate computational efficiency by measuring TFLOPS at inference.
@@ 4.1 Experimental Settings
Task fail, trackback * 1
@@
Baselines. Our method is compared with Tree-Planner (Hu et al., 2024), SELF-REFINE (Madaan et al., 2023), and a supervised finetuned planner. They are all reproduced using finetuned Llama 3 8B (Dubey et al., 2024) (the original Llama cannot achieve meaningful SR due to format errors). In addition, we consider several baseline methods that call the GPT-3.5 API, including ProgPrompt (Singh et al., 2023), Zero-shot Planner (Huang et al., 2022) and two self-refining planners, Local Replan (Raman et al., 2022; Guo et al., 2023) and Global Replan (Shinn et al., 2023). Their results are for reference only. See Section B.2 for details.
Implementation Details. Our implementation is consistent with the baseline methods by finetuning from Llama 3 8B (Dubey et al., 2024) on the VirtualHome-Env training set (paired with the equilibrium points). The number of finetuning epochs is set to 6, and the learning rate is 0.0002. The world model is finetuned on all planner interactions for 5 epochs using the same learning rate. A greedy LLM sampling strategy is used in later refinement steps until convergence. Moreover, we implement the KV cache to speed up inference. Further details are provided in Section B.3.
4.2 Main Results
In particular, Table 2 compares the efficacy of our training scheme, equilibrium sequence modeling, against alternative self-correction methods. It shows more than 11% improvement over the prompting-based self-refinement method SELF-REFINE (Madaan et al., 2023) and more than 12% improvement over the system 2-based Tree-Planner (Hu et al., 2024). The former stems from the effectiveness of our training objective, while the latter is due to the flexible self-correction mechanism without extensive manual design. Meanwhile, we maintain a simplistic supervised finetuning fashion similar to the compared methods, without intricate reinforcement learning. These validates the effectiveness of equilibrium sequence modeling in robot task planning.
4.3 Visualization
Figure 4 further compares our self-correction process with the baseline methods. As can be seen, our method is better at incorporating environmental feedback to improve the plan, while the baselines fail by simply repeating the previous plan, or by making only local adjustments that are insufficient. The rationale behind is that our method can flexibly take into account feedback through forward passes of LLMs, allowing arbitrary changes based on its knowledge.In contrast, the tree-based alternative requires backtracking in a tree, which is costly and does not fully exploit the verbalized knowledge in the feedback. More qualitative results are illustrated in Figures 10, 11, 12, 13 and 14 of the appendix.
Table 3: Effectiveness of different types of feedback. They are measured under the constraint of up to 10 rounds of internal or external feedback. Our trained planner is able to take into account various types of feedback to refine the plan.
@@ 4.3 Visualization
4.4 Ablation Study
This section validates the effectiveness of our method in performance and efficiency. See Appendix C for more results.
(a) Scaling of performance w.r.t. inference-time computation
(b) Comparison of performance over different planning horizons
@@ 4.4 Ablation Study
(a) Improving efficiency through reusing equilibrium solutions
(b) Dynamic compute allocation for more complex plans
Figure 6: Ablation study on computational efficiency. Our method reuses equilibrium solutions and dynamically allocates inference compute to improve performance-efficiency tradeoff. The latter is measured by the number of fixed-point iterations before convergence with mean and std.
Scaling of performance. Here, we follow Brown et al. (2024); Snell et al. (2025); Wu et al. (2025); Jaech et al. (2024); Guo et al. (2025) in considering the scaling w.r.t. inference computation. The results in Figure 5(a) show that our method achieves better performance-computation tradeoff along with leading scaling w.r.t. inference computation. Thus, more inference budget can be allocated to improve its performance. Furthermore, in Figure 5(b), we show that its performance advantage is largely due to better long-horizon planning capabilities, achieving more than twice the success rate of baselines on extremely long plans (length$>$20).
Computational efficiency. Although our planner training is slower than baselines (36h vs. 12h) due to the equilibrium solving process for synthesizing training pairs ($\approx$24h), it exhibits a competitive inference efficiency. For example, our method takes 16h to evaluate, while Tree-Planner takes 24h. This can be attributed to our design of reusing equilibrium in nested equilibrium solving, As illustrates in Figure 6(a), it accelerates the convergence, achieving better performance ($>$55%) with significantly fewer interactions. Furthermore, Figure 6(b) shows that our planner dynamically allocates compute for tasks of different complexity.
@@ 5 Conclusion
This work proposes an equilibrium model-based LLM planner that is capable of self-refining plans from external and internal feedback. Unlike existing self-refinement methods based on prompting or sophisticated reinforcement learning, our proposed equilibrium sequence modeling allows simple supervised training of the self-refining planners. Moreover, it also enables the planner to efficiently incorporate environmental feedback or a world model for closed-loop planning. We implement the proposed approach on the VirtualHome-Env benchmark, and the experimental results suggest that it can dynamically allocate inference-time computation to achieve state-of-the-art robot task planning performance.
Impact Statement
This paper presents a supervised learning framework for improving the closed-loop long-horizon capabilities of LLM agents, with the potential to complement the prevailing reinforcement learning-based or prompting-based frameworks. However, we note that any such improvement in planning capabilities should be treated with caution. For example, the LLM agent could autonomously create sub-goals that are threatening to humans, e.g., gaining more control. This calls for further research into LLM interpretability and safety.
@@ A.2 Training Deep Equilibrium Models
Its proof is given in Section A.4. Due to the challenge of exactly computing the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ in the above gradient, existing work often approximate it via the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). Recently, Fung et al. (2022); Geng et al. (2021a) propose to approximate the inverse Jacobian term by $A\approx I$, the former proving it under strong theoretical assumptions. In practice, dropping the inverse Jacobian/Hessian has been used extensively in one-step gradient (Bolte et al., 2023; Luketina et al., 2016; Finn et al., 2017; Liu et al., 2019; Garima et al., 2020) and shown to be effective on Transformer-based LLMs (Choe et al., 2023).
@@ A.3 Transformer-based Deep Equilibrium Models
Recently, following the prevalence of Diffusion Transformers (Peebles & Xie, 2023), deep equilibrium models are extended to image generation tasks. Geng et al. (2023) propose generative equilibrium Transformers consisting of two modules, one using Transformer as the basic unit $f_{\theta}$ in the equilibrium model. Their method yields advanced one-step image generation results. Bai & Melas-Kyriazi (2024) replace most of the intermediate Transformer blocks with an equilibrium model, thus significantly reducing the number of parameters and memory usage for training and inference.
@@ A.4 Proof of Implicit Function Theorem
Proof of Theorem 3.1. To derive the loss gradient w.r.t. $\theta$, we begin by differentiating the equilibrium condition $x^{*}=f_{\theta}(x^{*},c)$ with respect to $\theta$. Applying the chain rule, we have:
@@ Appendix B Experimental Settings
We detail the benchmark in Section B.1, the baselines in Section B.2, and our implementation details in Section B.3.
@@
Figure 8: Detailed statistics of VirtualHome-Env (Puig et al., 2018; Liao et al., 2019). It features (a) a large set of long-horizon plans with an average length of 10.8, and (b) 7 complex scenes containing more than 280 objects and more than 400 valid relations. For clarity, we exclude the CLOSE and FACING relations, which are redundant for most planning tasks.
@@ B.1 Benchmark
Figure 7 visualizes a few examples sampled from the VirtualHome-Env benchmark. In each example, the planner is placed in an environment that spans a few indoor rooms and is given a detailed description of the environment. The description is originally in the form of a scene graph with objects as nodes and spatial relationships as edges, but for simplicity we present the planner with only the object nodes. The planner is then asked to generate a semantic action sequence based on a short task description. For instance, after receiving an instruction “turn on TV …”, the robot agent must first walk to the table and grab the remote control, and then point at the TV to turn it on. As seen, these planning tasks usually involve a rather complex scene setup, and the ground truth action sequences are quite long. We provide more detailed statistics in Figure 8.
Compared to alternative embodied planning benchmarks, VirtualHome-Env features both long time horizons and closed-loop feedback. For example, ALFRED (Shridhar et al., 2020) and ReALFRED (Kim et al., 2024) are two common embodied instruction following benchmarks, but their plan lengths are relatively short and can be determined by a few templates, making them unsuitable for long-horizon planning. PlanBench (Valmeekam et al., 2023) and TravelPlanner (Xie et al., 2024) are recent benchmarks designed specifically for LLM planning, but they do not provide closed-loop feedback during execution, which is an essential element of robotic planning. Therefore, we adopt VirtualHome-Env during the experiments.
@@
    Program format feedback:“Your output does not conform to the required format”, indicating that the generated action sequence does not conform to the required format.
@@
    Invalid command feedback: “Your output has an invalid command: …”, indicating that the generated action sequence has an illegal command line.
@@
    Execution feedback: “Your output is executed incorrectly in the environment.”, indicating that the generated action sequence cannot be executed in the environment.
@@
    Task completion feedback: “You have not completed this task. The following objects and corresponding states do not meet the goals: … The following objects have wrong relative position: …”, indicating that the generated action sequence cannot complete the task, with more details about the task failure.
Table 4: Comparison of dataset protocols of VirtualHome (Puig et al., 2018). Since previous works have not released their dataset, we use the original VirtualHome-Env dataset (Liao et al., 2019) and perform our own partitioning.
Public #Tasks #Scenes Task Content Task Splits
Tree-Planner 35 4 Household Novel scene and task
LLM-MCTS 2000 4 Rearrangement Novel scene and task, Novel scene, Novel task, Seen scene and task
Ours $\checkmark$ 1360 7 Household Novel scene and task, Novel scene, Novel task
@@
Note that there are two alternative dataset protocols for VirtualHome, represented by LLM-MCTS (Zhao et al., 2023) and Tree-Planne (Hu et al., 2024), as summarized in Table 4. Specifically, LLM-MCTS uses the indoor scenes of VirtualHome and synthesizes its own dataset focusing on object rearrangement tasks. Tree-Planner uses a subset of VirtualHome-Env to evaluate for training-free methods. However, since neither works has released their detailed dataset, therefore we simply use the original VirtualHome-Env dataset and partition it accordingly in our experiments.
@@ B.2 Baselines
Our method is mainly compared with Tree-Planner (Hu et al., 2024) and SELF-REFINE (Madaan et al., 2023), both reproduced using Llama 3 8B Instruct (Dubey et al., 2024) in line with ours. The former traverses an action tree that is built by repeated plan sampling, while the latter relies on self-refinement. To reproduce them, we perform supervised finetuning of Llama 3 on the training split of VirtualHome-Env for the same number of epochs as our method, and then follow their original procedures for inference. For instance, Tree-Planner is reproduced with both settings $N\in\{25,50\}$ in action tree construction. The system prompts they use are similar to ours in Figure 9.
We also report the results summarized by Hu et al. (2024) for reference. They additionally considered ProgPrompt (Singh et al., 2023), Zero-shot Planner (Huang et al., 2022) and two self-refinement planners, Local Replan (Raman et al., 2022; Guo et al., 2023) and Global Replan (Shinn et al., 2023). Since these baselines were implemented by calling the GPT-3.5 API instead of finetuning Llama 3, we report them in the novel scene and task track for a relatively fair comparison. It is worth noting that they adopted a smaller subset of actions and feedback, and differed in the curation of partial observations of the environment. Therefore, their results are presented for reference only.
There are several robotic planning baselines that we have not compared due to large environmental differences. For example, LLM-MCTS (Zhao et al., 2023) is a representative tree-search (Yao et al., 2023a) based planner. It followed Watch-and-help (Puig et al., 2021) to generate a dataset of simple embodied tasks (mostly object rearrangement tasks), while our work considers a more complex set of planning tasks, see Table 4. Alternative planners based on symbolic scene graph (Zhu et al., 2021; Rana et al., 2023), code (Liang et al., 2023; Sun et al., 2023), or PDDL (McDermott, 2000; Liu et al., 2023; Guan et al., 2023) are less flexible and difficult to implement in our environment.